OpenStudy (anonymous):

Write an equation in slope-intercept form of the line with the given slope and point: slope = 3 and (1, -2). How do I solve this?

OpenStudy (anonymous):

y = mx + b -2 = 3(1) + b solve for b Once you have b, rewrite the equation y = 3 x + whatever you got

OpenStudy (anonymous):

Oh, I see now. Thank you so much!

OpenStudy (mathmale):

Alternatively, you could substitute the slope, m, and the point, (1, -2) into the point-slope form of the equation of a straight line, and then solve for y. I tend to favor this approach.

OpenStudy (anonymous):

I do, too...it's much easier...fewer steps

OpenStudy (anonymous):

y +2 = 3(x-1) y+2 = 3x+3 y = 3x+1 Is this what you got Maikel?